# Has your pretrained model improved? a multi-head posterior based approach

### 02 Jan 2024 v1

@@ 6 Conclusion
This work introduces a novel method for evaluating pretrained models. Instead of using costly and time-consuming fine-tuned downstream tasks for evaluation, we propose using the consistency between entity embeddings and their associated meta features as a performance metric. Our method has been effectively tested across various domains and datasets in relational datasets, Natural Language Processing and Computer Vision, providing a more efficient and equally rigorous alternative for pretrained model evaluation.
Appendix
@@
Appendix A Fixed regularization cross Model analysis of CLIP Embeddings on the BREEDS hierarchy
In this section we demonstrate supplementary results to the finding shown in Figure 10 regarding the dissonance between linear probe accuracies and log posterior values over the performance of some layers in the CLIP models. Figure 11 demonstrates the findings for entity 30, Figure 12 for living 17 and Figure 13 for non living 26.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 30 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
@@ Appendix A Fixed regularization cross Model analysis of CLIP Embeddings on the BREEDS hierarchy
Figure 11: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over entity 30 Breeds Dataset.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for living 17 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
@@
Figure 12: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over living 17 Breeds Dataset.
(a) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. Class covariance matrix assumes uncorrelated embedding dimensions, thereby just using the diagonal of the covariance matrix.
(b) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-3}$
(c) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for non living 26 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-6}$
(d) Layerwise MAP (Log. Posterior) value vs layerwise linear probe value for entity 13 breeds dataset. The regularization used to learn the class covariance matrix was $10^{-9}$
@@
Figure 13: Layerwise multi model analysis of log posterior performance with linear probe performance for a given regularization parameter over non living 26 Breeds Dataset.
Appendix B Analysis of CLIP Embeddings for ImageNet Subsets
In this section we continue the analysis on CLIP Image Encoder (Radford et al., 2021) embeddings on subsets of ImageNet based on superclass as defined by the Robustness library (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020). The experimental setup is explained next.
B.1 Datasets
To Test our approach we use 5 ImageNet(Deng et al., 2009) subsets using superclass labels based on WordNet(Miller, 1994) hierarchy from the robustness benchmarks (Engstrom et al., 2019a; b; Santurkar et al., 2019; 2020). The datasets are namely: big-12, geirhos-16, living-9, mixed-10, mixed-13. The dataset size statistics are included in Table 7 and the dataset label composition in Table 8.
Dataset Statistics of Robustness ImageNet Superclasses
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Training Samples 308494 40517 92632 77237 100233
Testing Samples 12005 1600 3603 3000 3903
Classes 12 16 9 10 13
Table 7: Dataset Statistics
Dataset
Class Composition
big-12
Dog (n02084071), Structure (n04341686), Bird (n01503061), Clothing (n03051540), Vehicle(n04576211), Reptile (n01661091), Carnivore (n02075296), Insect (n02159955), Instrument (n03800933), Food (n07555863), Furniture (n03405725), Primate (n02469914)
geirhos-16
Aircraft (n02686568), Bear (n02131653), Bicycle (n02834778), Bird (n01503061), Boat (n02858304), Bottle (n02876657), Car (n02958343), Cat (n02121808), Char (n03001627), Clock (n03046257), Dog (n02084071), Elephant (n02503517), Keyboard (n03614532), Knife (n03623556), Oven (n03862676), Truck (n04490091)
living-9
Dog (n02084071), Bird (n01503061), Arthropod (n01767661), Reptile (n01661091), Primate (n02469914), Fish (n02512053), Feline (n02120997), Bovid (n02401031), Amphibian (n01627424)
mixed-10
Dog (n02084071), Bird (n01503061), Insect (n02159955), Monkey (n02484322), Car (n02958343), Cat (n02120997), Truck (n04490091), Fruit (n13134947), Fungus (n12992868), Boat (n02858304)
mixed-13
Dog (n02084071), Bird (n01503061), Insect (n02159955), Furniture (n03405725), Fish (n02512053), Monkey (n02484322), Car (n02958343), Cat (n02120997), Truck (n04490091), Fruit (n13134947), Fungus (n12992868), Boat (n02858304), Computer (n03082979)
Table 8: ImageNet Superclass subsets
B.2 Experiments and Analysis
Given this setup we analyze the embeddings of the pre-trained CLIP Vision Transformers on the 2 data splits used to learn the parameters for the posterior distribution. For each network we compute the embeddings over all the datasets (train and test) shown in Table 7 and use the training set to learn a linear classifier in order to quantify the linear separability of CLIP embeddings and also learn the parameters of the posterior to assess the quality of embeddings. We plot the data and network wise posterior ($\log$) probability and the linear classifier accuracy and show the pearson correlation ($\rho$) between the 2 quantities for each respective network in Figure 14. 13(a) contain the results for the training set and the test set results are shown in 13(b).
(a) Train Posterior and linear probe plots
(b) Test Learned Posterior and linear probe plots
(c) Test Independent Posterior and linear probe plots
Figure 14: CLIP Embedding Posterior Analysis.
Training Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.219 -0.014 -0.064 -0.081 -0.102
Clustering Acc. Post. 95.336 99.721 98.517 98.191 97.852
Linear Probe Acc. 96.985 99.250 98.625 98.226 98.357
Table 9: ViT-B-16 Train Posterior, clustering performance and linear probe statistics
Training Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.271 -0.020 -0.109 -0.112 -0.137
Clustering Acc. Post. 94.209 99.595 97.471 97.477 97.105
Linear Probe Acc. 95.700 98.729 97.254 97.407 97.538
Table 10: ViT-B-32 Train Posterior, clustering performance and linear probe statistics
Training Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.125 -0.001 -0.012 -0.022 -0.029
Clustering Acc. Post. 97.356 99.975 99.732 99.512 99.388
Linear Probe Acc. 97.604 99.691 99.440 98.978 98.962
Table 11: ViT-L-14 Train Posterior, clustering performance and linear probe statistics
A more detailed presentation of results presented in 13(a) is shown in Table 9, Table 10 and Table 11. Further analyzing 13(a) for the training set we observe a strong correlation ($\rho$) between the $\log$ posterior probabilities and linear classifier accuracies. We also observe that for a given dataset a higher $\log$ posterior coincides with a higher linear classifier accuracy and a better clustering accuracy due to a higher posterior probability. We omitted the clustering accuracy plots for the posterior for brevity. But Table 9 - Table 17 contain those values.
Test Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Test Log. Post. -0.265 -0.225 -0.133 -0.168 -0.186
Clustering Acc. Post. 94.319 95.500 96.947 96.300 96.131
Linear Probe Test Acc. 97.143 99.125 98.279 98.200 98.411
Table 12: ViT-B-16 Test Posterior, clustering performance and linear probe statistics
Test Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.313 -0.254 -0.200 -0.207 -0.225
Clustering Acc. Post. 93.294 94.938 95.420 95.433 95.311
Linear Probe Acc. 94.494 97.688 97.197 97.833 97.771
Table 13: ViT-B-32 Test Posterior, clustering performance and linear probe statistics
Test Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.172 -0.180 -0.064 -0.082 -0.112
Clustering Acc. Post. 96.335 96.438 98.557 98.167 97.643
Linear Probe Acc. 97.943 99.375 99.223 99.100 99.154
Table 14: ViT-L-14 Test Posterior, clustering performance and linear probe statistics
Next, we continue the same analysis for 13(b) of the test set and its accompanying metrics in Table 12, Table 13 and Table 14. The first observation that carries over from our previous analysis on the training set is that for a given dataset, networks with higher posterior also have better accuracies, however as shown in 13(b) the correlation between those 2 sets of variables has reduced. Based on these observations we further investigate this by learning a posterior solely on the test data and comparing its generalization to the linear classifier and present the results in 13(c) and the corresponding metrics in Table 15, Table 16 and Table 17. Paying close attention to performance of ViT-B-16 and ViT-B-32 on the big-12 dataset we observe that while ViT-B-16 has more linearly separable embeddings across the training and test as shown by better linear probe accuracy in comparison to ViT-B-32. The $\log$ posterior and therefore the corresponding clustering accuracy of ViT-B-32 is higher than ViT-B-16, this presents a counter example to the earlier analysis. Further, looking at ViT-B-32 and ViT-L-14 we observe that $\log$ posterior for both networks is the same for all the datasets but the linear probe performance of those embeddings have variations. This non dependence between linear probe values and $\log$ posterior for perfectly clusterable data leads to an undefined Pearson Coefficient ($\rho$) as indicated by the value nan in 13(c).
Test (Ind.) Set Metrics - ViT-B-16
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. -0.0004 0.000 0.000 0.000 0.000
Clustering Acc. Post. 99.992 100.000 100.000 100.000 100.000
Linear Probe Acc. 97.143 99.125 98.279 98.200 98.411
Table 15: ViT-B-16 Independently learned Test Posterior, its clustering performance and original linear probe statistics
Test (Ind.) Set Metrics - ViT-B-32
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. 0.000 0.000 0.000 0.000 0.000
Clustering Acc. Post. 100.000 100.000 100.000 100.000 100.000
Linear Probe Acc. 94.494 97.688 97.197 97.833 97.771
Table 16: ViT-B-32 Independently learned Test Posterior, its clustering performance and original linear probe statistics
Test (Ind.) Set Metrics - ViT-L-14
Datasets big-12 geirhos-16 living-9 mixed-10 mixed-13
Train Log. Post. 0.000 0.000 0.000 0.000 0.000
Clustering Acc. Post. 100.000 100.000 100.000 100.000 100.000
Linear Probe Acc. 97.943 99.375 99.223 99.100 99.154
Table 17: ViT-L-14 Independently learned Test Posterior, its clustering performance and original linear probe statistics

### 14 Feb 2024 v4

@@ 6 Conclusion
This work introduces a novel method for evaluating pre-trained models. Instead of using costly and time-consuming fine-tuned downstream tasks for evaluation, we propose using the consistency between entity embeddings and their associated meta-features as a performance metric. Our method has been effectively tested across various domains and datasets in relational datasets, Natural Language Processing, and Computer Vision, providing a more efficient and equally rigorous alternative for pre-trained model evaluation.
Appendix A Appendix
@@
A.1 Fixed regularization cross Model analysis of CLIP Embeddings on the BREEDS hierarchy
In this section, we break the constituents of 6(a), Figure 8, Table 3 and Table 4 into individual plots comparing the behavior of ALP and linear probe accuracy for each CLIP model on a given dataset for varying regularization schemes. A complete per regularization breakdown for the 3 CLIP Models corresponding to entity-13 from 6(a) is provided in Figure 9 and Figure 10, for entity-30 in 6(b), the same is shown in Figure 11 and Figure 12. For living-17 and non-living-26 from 7(a) and 7(b), the analysis is shown in Figure 13, Figure 14 and Figure 15, Figure 16 respectively.
(a) Regularization - Uncorr.
(b) Regularization $10^{-3}$.
@@ A.1 Fixed regularization cross Model analysis of CLIP Embeddings on the BREEDS hierarchy
Figure 9: Regularization-wise analysis of layerwise ALP vs. linear probe performance for entity 13 breeds dataset.
(a) Regularization $10^{-6}$.
(b) Regularization $10^{-9}$.
@@
Figure 10: Regularization-wise analysis of layerwise ALP vs. linear probe performance for entity 13 breeds dataset.
(a) Regularization - Uncorr.
(b) Regularization $10^{-3}$.
@@
Figure 11: Regularization wise analysis of layerwise ALP vs. linear probe performance for entity 30 breeds dataset.
(a) Regularization $10^{-6}$.
(b) Regularization $10^{-9}$.
Figure 12: Regularization wise analysis of layerwise ALP vs. linear probe performance for entity 30 breeds dataset.
(a) Regularization - Uncorr.
(b) Regularization $10^{-3}$.
Figure 13: Regularization wise analysis of layerwise ALP vs. linear probe performance for living 17 breeds dataset.
(a) Regularization $10^{-6}$.
(b) Regularization $10^{-9}$.
Figure 14: Regularization wise analysis of layerwise ALP vs. linear probe performance for living 17 breeds dataset.
(a) Regularization - Uncorr.
(b) Regularization $10^{-3}$.
Figure 15: Regularization-wise analysis of layerwise ALP vs. linear probe performance for nonliving 26 breeds dataset.
(a) Regularization $10^{-6}$.
(b) Regularization $10^{-9}$.
Figure 16: Regularization-wise analysis of layerwise ALP vs. linear probe performance for nonliving 26 breeds dataset.